I don't have my book yet and my pretest question asks:
Find the inverse of: f(x)= x/4 (x over 4). I have no idea what inverse means pertaining to fractions. Thank you!!
Allison

Answer by Mike(39)

Here is what you need to do:
f(x) is the same as saying y. Therefore, f(x) = x/4 is the same as y = x/4. To find the inverse, simply switch y and x around. As such: x = y/4. Now all you have to do is solve for y by multiplying both sides by 4, giving you y = 4x, and that's the answer. The answer may also be expressed as follows: f^-1(x) = 4x. f^-1(x) means it is the inverse of the function.
Hope that helps.
Regards,
-Mike